# Scaling Depth Can Enable New Goal-Reaching Capabilities

### v2

@@ 4.5 Deeper Networks Enable Partial Experience Stitching
Figure 7: Testing the limits of scale. We extend the results from Figure 1 by scaling networks even further on the challenging Humanoid maze environments. We observe continued performance improvements with network depths of 256 and 1024 layers on Humanoid U-Maze. Note that for the 1024-layer networks, we observed the actor loss exploding at the onset of training, so we maintained the actor depth at 512 while using 1024-layer networks only for the two critic encoders.
4.6 Training Contrastive RL with 1000+ Layers
We next study whether further increasing depth beyond 64 layers further improves performance. We use the Humanoid maze tasks as these are both the most challenging environments in the benchmark and also seem to benefit from the deepest scaling. The results, shown in Figure 7, indicate that performance continues to substantially improve as network depth reaches 256 and 1024 layers in the Humanoid U-Maze environment. While we were unable to scale beyond 1024 layers due to computational constraints, we expect to see continued improvements with even greater depths, especially on the most challenging tasks.
Figure 8: Scaling network width vs. depth. Here, we reflect findings from previous works (Lee et al.,, 2024; Nauman et al., 2024b, ) which suggest that increasing network width can enhance performance. However, in contrast to prior work, our method is able to scale depth, yielding more impactful performance gains. For instance, in the Humanoid environment, raising the width to 4096 (depth=4) fails to match the performance achieved by simply doubling the depth to 8 (width=256). This comparative advantage of scaling depth seems more pronounced as the dimensionality of the observations increases.
4.7 Ablating Width vs. Depth
Past literature has shown that scaling network width can be effective  (Lee et al.,, 2024; Nauman et al., 2024b, ). In Figure 8, we find that scaling width is also helpful in our experiments: wider networks (2048 and 4096) consistently outperform narrower networks across three diverse environment types (depth held constant at 4). However, depth seems to be a more effective axis for scaling: simply doubling the depth to 8 (width held constant at 256) layers outperforms the widest networks on the Humanoid environment. The advantage of depth scaling is most pronounced in the Humanoid environment (observation dimension 268), followed by Ant Big Maze (dimension 29) and Arm Push Easy (dimension 17), suggesting that the comparative benefit may increase with higher observation dimensionality.
Note that the parameter count scales linearly with width but quadratically with depth. For comparison, a network with 4 MLP layers and 4096 hidden units has roughly 70M parameters, while one with a depth of 32 and 256 hidden units has only around 2M. Therefore, when operating under a fixed FLOP compute budget or specific memory constraints, depth scaling may be a more computationally efficient approach to improving network performance.
Figure 9: Critical depth and residual connections. Incrementally increasing the depth of shallow networks results in marginal performance gains (left). However, when the network depth reaches a critical threshold, performance improves dramatically (right). Residual connections are essential to stabilize the training of large networks in order to fully leverage the benefits of greater depth.
Figure 10: Deeper networks unlock batch size scaling. Scaling batch size has been an effective mechanism in many areas in ML (Chen et al.,, 2022; Zhang et al.,, 2024), but it has not demonstrated the same effectiveness in reinforcement learning (Obando-Ceron et al.,, 2023). Our findings indicate that, while scaling batch size provides only marginal benefits at the original network capacity, larger networks can effectively leverage batch size scaling to achieve further improvements.
4.8 Deep Networks Unlock Batch Size Scaling
Scaling batch size has been well-established in other areas of machine learning (Chen et al.,, 2022; Zhang et al.,, 2024). However, this approach has not translated as effectively to reinforcement learning (RL), and prior work has even reported negative impacts on value-based RL (Obando-Ceron et al.,, 2023). Indeed, in our experiments, simply increasing the batch size for the original CRL networks yields only marginal differences in performance (Figure 10, top left).
To test this hypothesis, we study the effect of increasing the batch size for networks of varying depths. As shown in Figure 10, scaling the batch size becomes effective as network depth grows. This finding offers evidence that by scaling network capacity, we may simultaneously unlock the benefits of larger batch size, potentially making it an important component in the broader pursuit of scaling self-supervised RL.
4.9 Ablating the Actor vs. Critic Networks
Figure 11: Actor vs. Critic. In the Arm Push Easy environment, scaling the critic is more effective, whereas in the Ant Big Maze, scaling the actor has a greater impact. In the Humanoid environment, scaling both actor and critic together is essential. Overall, the results suggest that scaling both actor and critic networks can complement each other to enhance performance.
To investigate the role of scaling in the actor and critic networks, Figure 11 presents the final performance for various combinations of actor and critic depths across three environments. Prior work (Nauman et al., 2024b, ; Lee et al.,, 2024) focuses on scaling the critic network, finding that scaling the actor degrades performance. In contrast, while we do find that scaling the critic is more impactful in two of the three environments (Humanoid, Arm Push Easy), our method benefits from scaling the actor network jointly, with one environment (Ant Big Maze) demonstrating actor scaling to be more impactful. These findings suggest that scaling both the actor and critic networks can play a complementary role in enhancing performance.
@@ 4.10 Does Depth Scaling Improve Offline Contrastive RL?
In preliminary experiments, we also evaluated depth scaling in the offline goal-conditioned setting using OGBench  (Park et al.,, 2024). We found no evidence that increasing the network depth of CRL improves performance in this offline setting. To further investigate this, we conducted ablations: (1) scaling critic depth while holding the actor at 4 or 8 layers, and (2) applying cold initialization to the final layers of the critic encoders (Zheng et al.,, 2024). In all cases, baseline depth 4 networks had the highest success. This finding hints that the reason why scaling depth is useful in the online setting might be partially because it aids in exploration.
Summary of key empirical findings: CRL is scalable to depths unattainable by other RL proprioceptive algorithms (1000+ layers), perhaps due to its self-supervised nature. Both width and depth are key factors influencing CRL’s performance, but depth achieves greater performance and better parameter-efficiency (similar performance for 50×50$\times$50 × smaller models). We observe signs of emergent behaviors in CRL with deep neural networks, such as humanoid learning to walk and navigate a maze. Scale unlocks learning difficult maze topologies. Batch size scaling occurs in CRL for deep networks. CRL benefits from both the actor and critic scale.
@@ Acknowledgments.
We gratefully acknowledge Galen Collier, Director of Researcher Engagement, for support through the NSF grant, as well as the staff at Princeton Research Computing for their invaluable assistance. We also thank Colin Lu for his discussions and contributions to this work. This research was partially supported by the National Science Centre, Poland (grant no. 2023/51/D/ST6/01609), and the Warsaw University of Technology through the Excellence Initiative: Research University (IDUB) program.
@@ Appendix A Additional Experiments
A.1 Python Environment Differences
In all plots presented in the paper, we used MJX 3.2.6 and Brax 0.10.1 to ensure a fair and consistent comparison. During development, we noticed discrepancies in physics behavior between the environment versions we employed (the CleanRL version of JaxGCRL) and the version recommended in a more recent commit of JaxGCRL (Bortkiewicz et al.,, 2024). Upon examination, the performance differences (shown in Figure 12) stem from a difference in versions in the MJX and Brax packages. Nonetheless, in both sets of MJX and Brax versions, performance scales monotonically with depth.
Figure 12: Scaling behavior for humanoid in two different python environments: MJX=3.2.3, Brax=0.10.5 and MJX=3.2.6, Brax=0.10.1 (ours) version of JaxGCRL. Scaling depth improves the performance significantly for both versions. However, in the environment we used, training requires fewer environment steps to reach a marginally better performance than in other Python environment.
Appendix B Technical Details
@@
Our experiments use JaxGCRL suite of GPU-accelerated environments and contrastive RL algorithm with hyperparameters reported in Table 2.
@@
batch_size 512 for the majority of experiments
action_repeat 1

### v3

@@ 4.5 Why Scaling Happens
Figure 12: Testing the limits of scale. We extend the results from Figure˜1 by scaling networks even further on the challenging Humanoid maze environments. We observe continued performance improvements with network depths of 256 and 1024 layers on Humanoid U-Maze. Note that for the 1024-layer networks, we observed the actor loss exploding at the onset of training, so we maintained the actor depth at 512 while using 1024-layer networks only for the two critic encoders.
@@ 4.6 Does Depth Scaling Improve Offline Contrastive RL?
In preliminary experiments, we evaluated depth scaling in the offline goal-conditioned setting using OGBench  (Park et al., , 2024). We found little evidence that increasing the network depth of CRL improves performance in this offline setting. To further investigate this, we conducted ablations: (1) scaling critic depth while holding the actor at 4 or 8 layers, and (2) applying cold initialization to the final layers of the critic encoders (Zheng et al., , 2024). In all cases, baseline depth 4 networks often had the highest success. A key direction for future work is to see if our method can be adapted to enable scaling in the offline setting.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgments.
We gratefully acknowledge Galen Collier, Director of Researcher Engagement, for support through the NSF grant, as well as the staff at Princeton Research Computing for their invaluable assistance. We also thank Colin Lu for his discussions and contributions to this work. This research was also partially supported by the National Science Centre, Poland (grant no. 2023/51/D/ST6/01609), and the Warsaw University of Technology through the Excellence Initiative: Research University (IDUB) program. Finally, we would also like to thank Jens Tuyls and Harshit Sikchi for providing helpful commends and feedback on the manuscript.
@@ Appendix A Additional Experiments
A.1 Scaled CRL Outperforms All Other Baselines on 8 out of 10 Environments
Figure 12: Scaled CRL (Ours) outperforms baselines CRL (original), SAC, SAC+HER, TD3+HER, GCSL, and GCBC in 8 out 10 environments.
In Figure 1, we demonstrated that increasing the depth of the CRL algorithm leads to significant performance improvements over the original CRL (see also Table 1). Here, we show that these gains translate to state-of-the-art results in online goal-conditioned RL, with Scaled CRL outperforming both standard TD-based methods such as SAC, SAC+HER, and TD3+HER, as well as self-supervised imitation-based approaches like GCBC and GCSL.
A.2 The CRL Algorithm is Key: Depth Scaling is Not Effective on Other Baselines
Figure 13: Depth scaling yields limited gains for SAC, SAC+HER, TD3+HER, GCSL, and GCBC.
Next, we investigate whether increasing network depth in the baseline algorithms yields similar performance improvements as observed in CRL. We find that SAC, SAC+HER, and TD3+HER do not benefit from depths beyond four layers, which is consistent with prior findings (Lee et al., , 2024; Nauman et al., 2024b, ). Additionally, GCSL and GCBC fail to achieve any meaningful performance on the Humanoid and Ant Big Maze tasks. Interestingly, we do observe one exception, as GCBC exhibits improved performance with increased depth in the Arm Push Easy environment.
Table 1: Increasing network depth (depth $D=4\rightarrow 64$) increases performance on CRL (Figure˜1). Scaling depth exhibits the greatest benefits on tasks with the largest observation dimension (Dim).
Task Dim D = 4 D = 64 Imprv.
Arm Binpick Hard 17 $38$ $\mathbf{\pm}4$ $219$ $\mathbf{\pm}15$ $5.7\times$
Arm Push Easy $308$ $\mathbf{\pm}33$ $762$ $\mathbf{\pm}30$ $2.5\times$
Arm Push Hard $171$ $\mathbf{\pm}11$ $410$ $\mathbf{\pm}13$ $2.4\times$
Ant U4-Maze 29 $11.4$ $\mathbf{\pm}4.1$ $286$ $\mathbf{\pm}36$ $25\times$
Ant U5-Maze $0.97$ $\mathbf{\pm}0.7$ $61$ $\mathbf{\pm}18$ $63\times$
Ant Big Maze $61$ $\mathbf{\pm}20$ $441$ $\mathbf{\pm}25$ $7.3\times$
Ant Hardest Maze $215$ $\mathbf{\pm}8$ $387$ $\mathbf{\pm}21$ $1.8\times$
Humanoid 268 $12.6$ $\mathbf{\pm}1.3$ $649$ $\mathbf{\pm}19$ $52\times$
Humanoid U-Maze $3.2$ $\mathbf{\pm}1.2$ $159$ $\mathbf{\pm}33$ $50\times$
Humanoid Big Maze $0.06$ $\mathbf{\pm}0.04$ $59$ $\mathbf{\pm}21$ $1051\times$
A.3 Additional Scaling Experiments: Offline GCBC, BC, and QRL
We further investigate several additional scaling experiments. As shown in Figure˜14, our approach successfully scales with depth in the offline GCBC setting on the antmaze-medium-stitch task from OGBench. We find that our the combination of layer normalization, residual connections, and Swish activations is critical, suggesting that our architectural choices may be applied to unlock depth scaling in other algorithms and settings. We also attempt to scale depth for behavioral cloning and the QRL (Wang et al., 2023a, ) algorithm—in both of these cases, however, we observe negative results.
Figure 14: Our approach successfully scales depth in offline GCBC on antmaze-medium-stitch (OGBench). In contrast, scaling depth for BC (antmaze-giant-navigate, expert SAC data) and for both online (FetchPush) and offline QRL (pointmaze-giant-stitch, OGBench) yield negative results.
A.4 Can Depth Scaling also be Effective for Quasimetric Architectures?
Prior work (Wang et al., 2023b, ; Liu et al., , 2023) has found that temporal distances satisfy an important invariance property, suggesting the use of quasimetric architectures when learning temporal distances. Our next experiment tests whether changing the architecture affects the scaling properties of self-supervised RL. Specifically, we use the CMD-1 algorithm (Myers et al., , 2024), which employs a backward NCE loss with MRN representations. The results indicate that scaling benefits are not limited to a single neural network parametrization. However, MRN’s poor performance on the Ant U5-Maze task suggests further innovation is needed for consistent scaling with quasimetric models.
Figure 15: Performance of depth scaling on CRL augmented with quasimetric architectures (CMD-1).
A.5 Additional Architectural Ablations: Layer Norm and Swish Activation
We conduct ablation experiments to validate the architectural choices of layer norm and swish activation. Figure˜16 shows that removing layer normalization performs significantly worse. Additionally, scaling with ReLU significantly hampers scalability. These results, along with Figure˜6 show that all of our architectural components—residual connections, layer norm, and swish activations—are jointly essential to unlocking the full performance of depth scaling.
Figure 16: (Left) Layer Norm is essential for scaling depth. (Right) Scaling with ReLU activations leads to worse performance compared to Swish activations.
A.6 Can We Integrate Novel Architectural Innovations from the Emerging RL Scaling Literature?
Recently, Simba-v2 proposed a new architecture for scalable RL. Its key innovation is the replacement of layer normalization with hyperspherical normalization, which projects network weights onto the unit-norm hypersphere after each gradient update. As shown, the same depth-scaling trends hold when adding hyperspherical normalization to our architecture, and it further improves the sample efficiency of depth scaling. This demonstrates that our method can naturally incorporate new architectural innovations emerging in the RL scaling literature.
Table 2: Integrating hyperspherical normalization in our architecture enhances the sample efficiency of depth scaling.
Steps to reach $\geq$200 success
Depth 4 16 32
With 50 42
Without 64 54
Steps to reach $\geq$400 success
Depth 4 16 32
With 62 48
Without 75 64
Steps to reach $\geq$600 success
Depth 4 16 32
With 77 67
Without 77
A.7 Residuals Norms in Deep Networks
Prior work has noted decreasing residual activation norms in deeper layers (Chang et al., , 2018). We investigate whether this pattern also holds in our setting. For the critic, the trend is generally evident, especially in very deep architectures (e.g., depth 256). The effect is not as pronounced in the actor.
Figure 17: L2 norms of residual activations in networks with depths of 32, 64, 128, and 256.
A.8 Scaling Depth for Offline Goal-conditioned RL
Figure 18: To evaluate the scalability of our method in the offline setting, we scaled model depth on OGBench (Park et al., , 2024). In two out of three environments, performance drastically declined as depth scaled from 4 to 64, while a slight improvement was seen on antmaze-medium-stitch-v0. Successfully adapting our method to scale offline GCRL is an important direction for future work.
Appendix B Experimental Details
B.1 Environment Setup and Hyperparameters
Figure 19: The scaling results of this paper are demonstrated on the JaxGCRL benchmark, showing that they replicate across a diverse range of locomotion, navigation, and manipulation tasks. These tasks are set in the online goal-conditioned setting where there are no auxiliary rewards or demonstrations. Figure taken from (Bortkiewicz et al., , 2024).
@@ B.1 Environment Setup and Hyperparameters
Our experiments use the JaxGCRL suite of GPU-accelerated environments, visualized in Figure˜19, and a contrastive RL algorithm with hyperparameters reported in Table˜7. In particular, we use 10 environments, namely: ant_big_maze, ant_hardest_maze, arm_binpick_hard, arm_push_easy, arm_push_hard, humanoid, humanoid_big_maze, humanoid_u_maze, ant_u4_maze, ant_u5_maze.
B.2 Python Environment Differences
In all plots presented in the paper, we used MJX 3.2.6 and Brax 0.10.1 to ensure a fair and consistent comparison. During development, we noticed discrepancies in physics behavior between the environment versions we employed (the CleanRL version of JaxGCRL) and the version recommended in a more recent commit of JaxGCRL (Bortkiewicz et al., , 2024). Upon examination, the performance differences (shown in Figure˜20) stem from a difference in versions in the MJX and Brax packages. Nonetheless, in both sets of MJX and Brax versions, performance scales monotonically with depth.
Figure 20: Scaling behavior for humanoid in two different python environments: MJX=3.2.3, Brax=0.10.5 and MJX=3.2.6, Brax=$0.1$ (ours) version of JaxGCRL. Scaling depth improves the performance significantly for both versions. In the environment we used, training requires fewer environment steps to reach a marginally better performance than in other Python environment.
B.3 Wall-clock Time of Our Approach
We report the wall-clock time of our approach in Table˜3. The table shows results for depths of 4, 8, 16, 32, and 64 across all ten environments, and for the Humanoid U-Maze environment, scaling up to 1024 layers. Overall, wall-clock time increases approximately linearly with depth beyond a certain point.
Table 3: Wall-clock time (in hours) for Depth 4, 8, 16, 32, and 64 across all 10 environments.
Environment Depth 4 Depth 8 Depth 16 Depth 32 Depth 64
Humanoid 1.48 $\pm$ 0.00 2.13 $\pm$ 0.01 3.40 $\pm$ 0.01 5.92 $\pm$ 0.01 10.99 $\pm$ 0.01
Ant Big Maze 2.12 $\pm$ 0.00 2.77 $\pm$ 0.00 4.04 $\pm$ 0.01 6.57 $\pm$ 0.02 11.66 $\pm$ 0.03
Ant U4-Maze 1.98 $\pm$ 0.27 2.54 $\pm$ 0.01 3.81 $\pm$ 0.01 6.35 $\pm$ 0.01 11.43 $\pm$ 0.03
Ant U5-Maze 9.46 $\pm$ 1.75 10.99 $\pm$ 0.02 16.09 $\pm$ 0.01 31.49 $\pm$ 0.34 46.40 $\pm$ 0.12
Ant Hardest Maze 5.11 $\pm$ 0.00 6.39 $\pm$ 0.00 8.94 $\pm$ 0.01 13.97 $\pm$ 0.01 23.96 $\pm$ 0.06
Arm Push Easy 9.97 $\pm$ 1.03 11.02 $\pm$ 1.29 12.20 $\pm$ 1.43 14.94 $\pm$ 1.96 19.52 $\pm$ 1.97
Arm Push Hard 9.74 $\pm$ 1.05 10.55 $\pm$ 1.20 11.98 $\pm$ 1.49 14.40 $\pm$ 1.64 18.53 $\pm$ 0.06
Arm Binpick Hard 18.41 $\pm$ 2.16 17.48 $\pm$ 1.88 19.47 $\pm$ 0.05 21.91 $\pm$ 1.93 29.64 $\pm$ 6.10
Humanoid U-Maze 8.72 $\pm$ 0.01 11.29 $\pm$ 0.01 16.36 $\pm$ 0.03 26.48 $\pm$ 0.05 46.74 $\pm$ 0.04
Humanoid Big Maze 12.45 $\pm$ 0.02 15.02 $\pm$ 0.01 20.34 $\pm$ 0.01 30.61 $\pm$ 0.05 50.33 $\pm$ 0.05
Table 4: Total wall-clock time (in hours) for training from Depth 4 up to Depth 1024 in the Humanoid U-Maze environment.
Depth Time (h)
4 3.23 $\pm$ 0.001
8 4.19 $\pm$ 0.003
16 6.07 $\pm$ 0.003
32 9.83 $\pm$ 0.006
64 17.33 $\pm$ 0.003
128 32.67 $\pm$ 0.124
256 73.83 $\pm$ 2.364
512 120.88 $\pm$ 2.177
1024 134.15 $\pm$ 0.081
B.4 Wall-clock Time: Comparison to Baselines
Since the baselines use standard sized networks, naturally our scaled approach incurs higher raw wall-clock time per environment step (Table˜5). However, a more practical metric is the time required to reach a given performance level. As shown in Table˜6, our approach outperforms the strongest baseline, SAC, in 7 of 10 environments while requiring less wall-clock time.
Table 5: Wall-clock training time comparison of our method vs. baselines across all 10 environments.
Environment Scaled CRL SAC SAC+HER TD3 GCSL GCBC
Humanoid 11.0 $\pm$ 0.0 0.5 $\pm$ 0.0 0.6 $\pm$ 0.0 0.8 $\pm$ 0.0 0.4 $\pm$ 0.0 0.6 $\pm$ 0.0
Ant Big Maze 11.7 $\pm$ 0.0 1.6 $\pm$ 0.0 1.6 $\pm$ 0.0 1.7 $\pm$ 0.0 1.5 $\pm$ 0.3 1.4 $\pm$ 0.1
Ant U4-Maze 11.4 $\pm$ 0.0 1.2 $\pm$ 0.0 1.3 $\pm$ 0.0 1.3 $\pm$ 0.0 0.7 $\pm$ 0.0 1.1 $\pm$ 0.1
Ant U5-Maze 46.4 $\pm$ 0.1 5.7 $\pm$ 0.0 6.1 $\pm$ 0.0 6.2 $\pm$ 0.0 2.8 $\pm$ 0.1 5.6 $\pm$ 0.5
Ant Hardest Maze 24.0 $\pm$ 0.0 4.3 $\pm$ 0.0 4.5 $\pm$ 0.0 5.0 $\pm$ 0.0 2.1 $\pm$ 0.6 4.4 $\pm$ 0.5
Arm Push Easy 19.5 $\pm$ 0.6 8.3 $\pm$ 0.0 8.5 $\pm$ 0.0 8.4 $\pm$ 0.0 6.4 $\pm$ 0.1 8.3 $\pm$ 0.3
Arm Push Hard 18.5 $\pm$ 0.0 8.5 $\pm$ 0.0 8.6 $\pm$ 0.0 8.3 $\pm$ 0.1 5.2 $\pm$ 0.3 7.4 $\pm$ 0.5
Arm Binpick Hard 29.6 $\pm$ 1.3 20.7 $\pm$ 0.1 20.7 $\pm$ 0.0 18.4 $\pm$ 0.3 8.0 $\pm$ 0.9 16.2 $\pm$ 0.4
Humanoid U-Maze 46.7 $\pm$ 0.0 3.0 $\pm$ 0.0 3.5 $\pm$ 0.0 5.4 $\pm$ 0.0 3.1 $\pm$ 0.1 7.2 $\pm$ 0.8
Humanoid Big Maze 50.3 $\pm$ 0.0 8.6 $\pm$ 0.0 9.3 $\pm$ 0.0 7.5 $\pm$ 1.1 5.1 $\pm$ 0.0 11.4 $\pm$ 1.9
Table 6: Wall-clock time (in hours) for our approach to surpass SAC’s final performance. As shown, our approach surpasses SAC performance in less wall-clock time in 7 out of 10 environments. The N/A* entries are because in those environments, scaled CRL doesn’t outperform SAC.
Environment SAC Scaled CRL (Depth 64)
Humanoid 0.46 6.37
Ant Big Maze 1.55 0.00
Ant U4-Maze 1.16 0.00
Ant U5-Maze 5.73 0.00
Ant Hardest Maze 4.33 0.45
Arm Push Easy 8.32 1.91
Arm Push Hard 8.50 6.65
Arm Binpick Hard 20.70 4.43
Humanoid U-Maze 3.04 N/A*
Humanoid Big Maze 8.55 N/A*
@@ B.4 Wall-clock Time: Comparison to Baselines
batch_size 512